# StreamVoice: Streamable Context-Aware Language Modeling for Real-time Zero-Shot Voice Conversion

### v1

@@ 3.3 Context-aware Enhancement
Due to the disadvantage of the causality in the streaming framework, streaming models face missing future reception and potential performance degradation compared to the non-streaming model, while the low-quality semantic input from the streaming ASR, as we mentioned in Section 1, makes the streaming VC for high-quality conversion more challenging. To make the proposed streamable LM-based VC model perform conversion with high quality, a context-aware enhancement method is proposed, which can alleviate incomplete contextual information arising from the semantic input and the absence of future information. Specifically, we introduce context-masked autoregressive prediction in the LM to enhance the capture of historical context from the given semantic input, and meanwhile, a teacher-guided context foresight is proposed to ensure the model can imagine the future context based on that of its historical context.
3.3.1 Context-masked Autoregressive Prediction
@@
As shown in the left of Fig. 3, the LM is achieved by the multi-layer LLaMA with unidirectional attention. To enhance contextual awareness from the given semantic input, semantic masking is introduced in the LM to encourage acoustic prediction from the corrupted semantic. Specifically, within a sequence of semantic tokens $\mathbf{s}=\{s_{1},s_{2},...s_{T}\}$, we randomly select several indices as start indices at a ratio $r$, and spans of $l$ steps are masked by [M]. After masking, LM takes the corrupted semantic feature ${}^{m}\mathbf{s}$ as input and performs autoregression. With this method, an information bottleneck is also implicitly created in the semantic feature to reduce speaker information. Moreover, during training, we do not explicitly use a speech clip as the speaker prompt. Instead, LM leverages the previous sequence $\{\mathbf{s}_{1:t-1},\mathbf{a}_{1:t-1},s_{t}\}$ as prompts to autoregressively generate hidden representation $h_{t}$ for further acoustic prediction. Notably, during training, when the input of the current step is $a_{t}$, the corresponding output is skipped and does not involve further steps.
3.3.2 Teacher-guided Context Foresight
@@
As previously discussed, the absence of future information resulting in the loss of contextual information leads to a decline in the conversion performance. Inspired by the effective representation learning exhibited by autoregressive predictive coding Chung et al. (2019) (APC), we introduce teacher-guided context foresight guided by a non-streaming ASR to enhance the autoregression output, as presented in the right of Fig. 3. This allows the model to learn a context vector containing envisioned future information. Specifically, the context representation $\mathbf{c}$ is first derived by linear prediction from the hidden features $\mathbf{h}$, which is generated by the LM through historical context. Subsequently, this $c_{t}$ is encouraged to discover more general future context information by minimizing the L2 distance not only with $k$ semantic features from future time steps ${\overline{s}_{t+1},...,\overline{s}_{t+k}}$ but also with the current semantic feature $\overline{s_{t}}$. This dual minimization approach contributes to precise content delivery and enhances the model’s ability to forecast future context. The loss can be summarized as
@@
Furthermore, since the semantic feature $\{\mathbf{s},\mathbf{\overline{s}}\}$ from ASR still may contain speaker-related information. To further ensure the speech decoupling, the bottleneck regulator Qian et al. (2019), which squeezes out speaker information by reducing dimension size with a linear layer, is applied in $\mathbf{s}$ and $\mathbf{c}$.
@@ 3.4 Training & Inference Procedure
3.4.1 Training
During the training of StreamVoice, the context-enhanced language model and acoustic predictor are trained together. Both the autoregressive processes in these two parts are performed with a teacher-forcing strategy. The total loss can be described as $\mathcal{L}_{total}=\mathcal{L}_{TF}+\mathcal{L}_{Cont}$ for using continuous codec feature or $\mathcal{L}_{total}=\mathcal{L}_{TF}+\mathcal{L}_{Disc}$ for discrete version.
3.4.2 Streaming Inference
In inference, we use the semantic and acoustic features from a short speech clip of the target speaker as the speaker prompt. Since this clip is randomly selected, which may contain unfinished pronunciation at the end of the clip, we pad a silence clip after the speaker recording before the conversion process to prevent the unexpected continuation. With this prompt, StreamVoice can stream convert the source speech. In discrete projection, we use greedy decoding to choose the codec token with the highest probability. Besides, to ensure the real-time streaming inference of StreamVoice, we employ a commonly used key-value cache in LM to reduce redundant calculations. In our pipeline, since the beginning and end of the source speech can be determined by ASR or voice activity detection (VAD), we don’t employ techniques, such as window attention or slide attention, to handle the input. It is noteworthy that these techniques can be easily integrated into our framework, providing flexibility for future extensions.
Table 1: Zero-shot performance (unseen speakers)
Table 2: In-datatset performance (seen speakers)
@@ 3.4.2 Streaming Inference
Method Chunk (ms) Quality Similarity
GT - - 3.65 6.29 - 0.729
Non-streaming Topline
NS-VC - 3.85$\pm$0.09 3.71 8.39 3.92$\pm$0.08 0.744
Streaming Model
IBF-VC 160 3.71$\pm$0.09 3.48 9.52 3.67$\pm$0.10 0.687
DualVC2 160 3.80$\pm$0.10 3.57 10.2 3.85$\pm$0.09 0.703
StreamVoice 80 3.82$\pm$0.09 3.50 10.0 3.82$\pm$0.10 0.694
$\quad$++ Tuning 80 3.78$\pm$0.08 3.52 10.4 3.87$\pm$0.10 0.714
Table 2: In-datatset performance (seen speakers)
4 Experiments
We first present the experimental setup in our work. Next, we provide subjective and objective evaluations and the ablation study conducted on StreamVoice. A detailed analysis of dependency in the streaming pipeline is also provided.
4.1 Experimental Setup
4.1.1 Corpus
A mixed dataset comprising 1,500 hours of Aishell3 Shi et al. (2020) and an internal Chinese dataset are used to train StreamVoice and Audiodec Wu et al. (2023). To extract semantic features, we incorporate a streaming ASR Fast-U2++ Liang et al. (2023), which is implemented by WeNet and trained on WenetSpeech Zhang et al. (2022). For zero-shot testing, a set of 400 testing pairs is selected from DIDISpeech Guo et al. (2021) and EMIME Wester (2010), each with a source and target speaker utterance. For evaluation on seen speakers, eight speakers from Aishell3 are selected to form 160 conversion pairs.
4.1.2 Implement Details
We use open-sourced code [^2] https://github.com/facebookresearch/AudioDec of Audiodec, which has 4 quantizer layers with a 1024 codebook size and 64 codebook dimension, representing a 24kHz waveform in 20ms frame length. The Fast-U2++ uses an 80ms chunk size to perform streaming inference and compresses a 16kHz waveform into a semantic feature with a 40ms frame length. For StreamVoice, we employ the LLaMA Touvron et al. (2023) architecture for context-enhanced LM, with 6 layers and 8 heads. The hidden and intermediate sizes are 1024 and 4096. We use the officially released code [^3] https://github.com/b04901014/MQTTS to implement the acoustic predictor, which uses a layer Transformer decoder with a hidden size 256, feed-forward hidden size 1024, and 4 heads. In semantic masking, mask ratio $r$ ranges from $0.01$ to $0.02$, and span $l$ is set to 10. The foresight step $k$ is set to 4. The bottleneck regulator compresses feature dimensions by 6 times. During training, the max training length is set to 12s. StreamVoice is trained using 8 V100 GPUs with a batch size of 7 per GPU for 700k steps. We use the AdamW optimizer with a learning rate of $5\times 10^{-4}$. Exponential decay updates the learning rate after each epoch, using a decay ratio of 0.986.
4.1.3 Evaluation metrics
[…]
@@ 4.2.1 Zero-shot evaluation
To evaluate the zero-shot VC performance, one recent LM-based zero-shot VC system, LM-VC Wang et al. (2023c), is selected as the topline system. Besides, a variant of StreamVoice, referred to as NS-StreamVoice, using a non-streaming ASR for semantic extraction, is also compared. We implement the proposed system StreamVoice integration discrete projection, while C-StreamVoice also involves the evaluation since speech can represented in continuous form by codec model. Table. 3.4.2 presents both subjective and objective results. Compared with the non-streaming topline LM-VC, our proposed StreamVoice can achieve close results regarding subjective NMOS and SMOS, while a performance gap still exists. Similar results are also observed in objective results. The non-streaming StreamVoice even surpasses the topline model in certain aspects, indicating the effectiveness of our well-designed streamable architecture for zero-shot VC. As can be observed, C-StreamVoice exhibits inferior performance compared to the discrete version, which can contribute to the reported over-smoothness in speech generation Ren et al. (2022) and the mismatch between the ground-truth and predicted features.
Furthermore, as illustrated in Table. 3, the RTF of the entire pipeline is below 1, which meets the real-time requirement. Consisting of chunk-waiting latency (80ms) and model inference latency, the overall pipeline latency is 124.3 ms. If using a V100 GPU, StreamVoice can obtain an RTF of 0.56 and the overall latency reaches 137.2ms. Importantly, unlike previous streaming VC, our VC model is entirely causal without any future look-ahead, highlighting its powerful modeling capability. These results demonstrate that StreamVoice can achieve high-quality zero-shot VC in streaming scenarios.
Table 3: Speed of StreamVoice tested on a single A100 80G GPU. Latency is obtained by multiplying the RTF by the 80ms chunk size.
@@ 4.2.2 In-dataset Evaluation
To get further insight into StreamVoice, we conducted an in-dataset evaluation on eight seen speakers, as shown in Table. 3.4.2. A non-streaming VC system Tian et al. (2020) achieving any2many VC, is selected, referred to as NS-VC. Also, IBF-VC Chen et al. (2022) and DualVC2 Ning et al. (2023) are recently proposed streaming models for any2many VC. As observed, a performance gap exists between the strong non-streaming topline and streaming models. Among the streaming models, StreamVoice, designed for the zero-shot scenario, delivers similar results to systems designed for in-dataset speakers. Even though StreamVoice uses a smaller chunk size of 80ms in streaming ASR which achieves lower ASR performance, indicating the good conversion ability of StreamVoice. With available utterances of target speakers, fine-tuning the StreamVoice yields superior performance. This indicates our system can be easily applied to various scenarios with or without the utterances of target speakers.
Table 4: Results of ablation studies.
@@ 4.3 Ablation study
To further validate the effectiveness of StreamVoice, we conducted several ablations studies on teacher-guided context foresight, semantic masking, and the bottleneck regulator. Table 4 presents the results.
In w/o teacher-guided context foresight, we discard the prediction of current and future semantic information, forming two ablations w/o $\mathcal{L}_{TF}(\overline{s}_{t})$ and w/o $\mathcal{L}_{TF}(\overline{s}_{t+1:t+k})$, respectively. As can be seen, a noticeable decrease occurs in all evaluation metrics when the $\mathcal{L}_{TF}(\overline{s}_{t+1:t+k})$ is discarded, especially in WVMOS and CER. This indicates that this foresight improves performance in capturing the linguistic content. But when only integrating context from future semantics, the model w/o $\mathcal{L}_{TF}(\overline{s}_{t})$ faces severe performance loss. It shows that only using future information interferes with delivering current linguistic content.
In w/o semantic masking, the semantic masking does not apply to the semantic feature during the training of StreamVoice. We observe a performance decrease in all evaluation metrics when the masking is discarded. This indicates that StreamVoice, trained with semantic masking, effectively enhances contextual learning from the preceding input while improving speaker timbre capturing.
We also evaluated the effect of the bottleneck regulator by dropping this usage, referred to as w/o bottleneck regulator. The results show that the integration of the bottleneck regulator is effective in preventing the source speaker information contained in the semantic feature from leaking into the converted speech, with little effect on the speech quality.
Table 5: Analysis of dependency on ASR. The results represent the performance of StreamVoice integrating different ASR.
@@ 4.4 Discussion: Dependency Analysis
As shown in Fig. 2, in the streaming pipeline, additional ASR and codec are needed to extract semantic and acoustic information. In this section, we will explore the dependency relations between the selection of ASR and codec and the performance of StreamVoice.
4.4.1 ASR
To investigate the impact of ASR on StreamVoice, three representative ASR systems, including non-streaming ASR 5, widely used CTC-based streaming ASR Moritz et al. (2019), and the recently proposed streaming Fast-U2++ Liang et al. (2023), are selected to perform semantic extraction. As can be seen in Table 5, StreamVoice using semantic features of non-streaming ASR outperforms those using streaming ASR. This discrepancy may be attributed to the inherent performance gap between non-streaming and streaming ASR models, resulting in different semantic extraction abilities. Besides, without future look-ahead in StreamVoice, using semantic features from Moritz et al. (2019) cannot achieve reasonable conversion, while we introduce a 160ms future look-ahead in StreamVoice, i.e., modeling $p(a_{t}|\mathbf{a}_{1:t-1},\mathbf{s}_{1:t+m},t)$ with $m$ future look-ahead, yield good conversion results. This issue may arise from delayed CTC spike distributions and token emission latency existing in streaming ASR Liang et al. (2023), leading to semantic information shifting. Benefiting from the low emission latency of Fast-U2++, StreamVoice can perform conversion without future look-ahead. With a longer chunk size employed in Fast-U2++, StreamVoice can obtain better results, while reaching a larger latency of 270ms. A trade-off still exists between performance and speed within the streaming pipeline.
Table 6: Analysis of dependency on Audiodec with various bitrate.
@@ 4.4.1 ASR
4.4.2 Codec
In StreamVoice, we employ a low-latency streaming codec Audiodec Wu et al. (2023). As presented in Table. 6, we validate the performance of StreamVoice using codecs with different bitrates, including 2kbps and 8kbps, where higher bitrate codecs achieve superior reconstruction quality to lower bitrate ones. The 2kbps Audiodec utilizes 4 layers of quantization and represents audio with a frame length of 20ms, while the 8kbps Audiodec employs 8 layers with a frame length of 10ms. Using the configuration of StreamVoice mentioned in Section 4.1.2, the results in different bitrates of codec models show no obvious differences. When increasing the number of transformer layers in the codec predictor, forming Large w/ 8kbps Audiodec, conversion performance using 8kbps codec improves noticeably, but resulting in slower inference. This result shows that the design of StreamVoice depends on the codec configuration, affecting both conversion quality and inference speed.
4.5 Conclusions
In this paper, we introduce StreamVoice, a novel LM-based zero-shot VC system designed for streaming scenarios. Specifically, StreamVoice employs a streamlined, single-stage framework that encompasses a context-aware LM and an acoustic predictor. The casual design of the model’s input and structure ensures compliance with streaming behavior. To address performance degradation caused by missing complete contextual information in streaming scenarios, context-aware LM adopts teacher-guided context foresight to make the model have the ability to forecast the current and future information given by a teacher. Besides, semantic masking is introduced in LM to enhance context learning from historical input and facilitate better disentanglement. Finally, an acoustic predictor collaborates with the LM to generate the target speech. Experiments demonstrate that StreamVoice achieves streaming zero-shot VC while maintaining performance comparable to non-streaming VC systems.
Limitations and future work. We have to point out that StreamVoice still has limitations. In our configuration, StreamVoice needs a GPU, such as V100 and A100, to achieve real-time streaming inference. The design of streaming VC heavily relies on the ASR and codec as mentioned in Section 4.4. Besides, StreamVoice also faces the out-of-domain problem, which causes performance degradation for utterances with accents, strong emotions, or unseen recording environments. Our future work will first use more training data to explore the modeling ability of StreamVoice. Also, we will focus on optimizing our streaming pipeline, such as high fidelity codec with low bitrate and unified streaming model.
Ethics Statement

### v5

@@ 3.3 Context-aware Enhancement
Due to the disadvantage of the causality in the streaming framework, streaming models face missing future reception and potential performance degradation compared to the non-streaming model, while the low-quality semantic input from the streaming ASR, as we mentioned in Section 1, makes achieving high-quality conversion more challenging. To address these issues, a context-aware enhancement method is proposed, which can alleviate incomplete contextual information arising from the semantic input and the absence of future information. Specifically, we introduce context-masked autoregressive prediction in the LM to enhance the capture of historical context from the given semantic input. Meanwhile, a teacher-guided context foresight is proposed to ensure the model can imagine the future context based on that of its historical context.
@@
Context-masked Autoregressive Prediction. As shown in the left of Fig. 3, the LM is achieved by the multi-layer Transformer with unidirectional attention, following the implementation of LLaMA Touvron et al. (2023). To enhance contextual awareness from the given semantic input, semantic masking is introduced in the LM to encourage acoustic prediction from the corrupted semantic. Specifically, within a sequence of semantic tokens $\mathbf{s}=\{s_{1},s_{2},...s_{T}\}$, we randomly select several indices as start indices at a ratio $r$, and spans of $l$ steps are masked by [M]. After masking, LM takes the corrupted semantic feature ${}^{m}\mathbf{s}$ as input and performs autoregression. With this method, an information bottleneck is also implicitly created in the semantic feature to reduce speaker information. Moreover, during training, we do not explicitly use a speech clip as the speaker prompt. Instead, LM leverages the previous sequence $\{\mathbf{s}_{1:t-1},\mathbf{a}_{1:t-1},s_{t}\}$ as prompts to autoregressively generate hidden representation $h_{t}$ for further acoustic prediction. Notably, when the current input is $a_{t}$, the corresponding output is skipped and does not involve further steps.
@@
Teacher-guided Context Foresight. As previously discussed, the absence of future information resulting in the loss of contextual information leads to a decline in the conversion performance. Inspired by the effective representation learning exhibited by autoregressive predictive coding Chung et al. (2019) (APC), we introduce teacher-guided context foresight guided by a non-streaming ASR to enhance the autoregression output, as presented in the right of Fig. 3. This allows the model to learn a context vector containing envisioned future information. Specifically, the context representation $\mathbf{c}$ is first derived by linear prediction from the hidden features $\mathbf{h}$, which is generated by the LM through historical context. Subsequently, this $c_{t}$ is encouraged to discover more general future context information by minimizing the L2 distance not only with $k$ semantic features from future time steps ${\overline{s}_{t+1},...,\overline{s}_{t+k}}$ but also with the current semantic $\overline{s_{t}}$. This dual minimization approach contributes to precise content delivery and enhances the ability to forecast future context. The loss can be summarized as
@@
Furthermore, since the semantic feature $\{\mathbf{s},\mathbf{\overline{s}}\}$ still may contain speaker-related information. To further ensure the speech decoupling, the bottleneck regulator Qian et al. (2019), which squeezes out speaker information by reducing dimension size with a linear layer, is applied in $\mathbf{s}$ and $\mathbf{c}$.
@@ 3.4 Training & Inference Procedure
Training. During training, the context-enhanced language model and acoustic predictor are trained together. The total loss can be described as $\mathcal{L}_{total}=\mathcal{L}_{TF}+\mathcal{L}_{Cont}$ for continuous codec or $\mathcal{L}_{total}=\mathcal{L}_{TF}+\mathcal{L}_{Disc}$ for discrete version.
Streaming Inference. We use the semantic and acoustic features from a short speech clip of the target speaker as the speaker prompt. Since this clip is randomly selected, which may contain unfinished pronunciation at the end of the clip, we pad a silence clip after the speaker recording before the conversion process to prevent the unexpected continuation. With this prompt, StreamVoice can stream convert the source speech. In discrete projection, we use greedy decoding to choose the codec token with the highest probability. Besides, to ensure the real-time streaming inference of StreamVoice, we employ the key-value cache in LM to reduce redundant calculations. In practice, since the beginning and end of the source speech can be determined by ASR or voice activity detection (VAD), we don’t employ techniques, such as window attention or slide attention, to handle the input. StreamVoice’s performance decreases when the long input exceeds the maximum training length. Notably, these techniques can be easily integrated into our framework, providing flexibility for future extensions.
4 Experiments
4.1 Experimental Setup
Corpus. A mixed dataset comprising 1,500 hours of Aishell3 Shi et al. (2021) and an internal Chinese dataset are used to train StreamVoice and Audiodec Wu et al. (2023). The internal dataset contains recordings from 2679 Chinese speakers, while we use utterances from 200 speakers in Aishell3. To extract semantic features, we incorporate a streaming ASR Fast-U2++ Liang et al. (2023), which is implemented by WeNet Yao et al. (2021) and trained on WenetSpeech Zhang et al. (2022). For zero-shot testing, a set of 400 testing pairs is selected from DIDISpeech Guo et al. (2021) and EMIME Wester (2010), each with a source and target speaker utterance. For evaluation of seen speakers, eight speakers from Aishell3 are selected to form 160 conversion pairs. And 3s speech utterance is used as a speaker prompt in inference. The duration of testing utterances is between 3s and 7s.
Implement Details. We use open-sourced code [^1] https://github.com/facebookresearch/AudioDec of Audiodec, which has 4 quantizer layers with a 1024 codebook size and 64 codebook dimension, representing a 24kHz waveform in 20ms frame length. The Fast-U2++ uses an 80ms chunk size to perform streaming inference and compresses a 16kHz waveform into a semantic feature with a 40ms frame length. StreamVoice contains 101M parameters. For context-enhanced LM, we employ the variant of Transformer, LLaMA Touvron et al. (2023), with 6 layers and 8 heads. The hidden and intermediate sizes are 1024 and 4096. We use the officially released code [^2] https://github.com/b04901014/MQTTS to implement the acoustic predictor, which uses a layer Transformer decoder with a hidden size 256, feed-forward hidden size 1024, and 4 heads. In semantic masking, mask ratio $r$ ranges from $0.01$ to $0.02$, and span $l$ is set to 10. The foresight step $k$ is set to 4. The bottleneck regulator compresses feature dimensions by 6 times. During training, the max training length is set to 12s. StreamVoice is trained using 8 V100 GPUs with a batch size of 7 utterances per GPU for 700k steps. We use the AdamW optimizer with a learning rate of $5\times 10^{-4}$. Exponential decay updates the learning rate after each epoch, using a decay ratio of 0.986.
[…]
@@ 4.1 Experimental Setup
Table 1: Zero-shot performance (unseen speakers)
@@ 4.2.1 Zero-shot Evaluation
To evaluate the zero-shot VC performance, one recent LM-based zero-shot VC system, LM-VC Wang et al. (2023c), is selected as the topline system. Besides, a variant of StreamVoice, referred to as NS-StreamVoice, using a non-streaming ASR for semantic extraction, is also compared. We implement the proposed system StreamVoice integration discrete projection, while C-StreamVoice also involves the evaluation since speech can represented in continuous form by codec model. Table. 1 presents both subjective and objective results. Compared with the non-streaming topline LM-VC, our proposed StreamVoice can achieve close results regarding subjective NMOS and SMOS, while a performance gap still exists. Similar results are also observed in objective results. The non-streaming StreamVoice even surpasses the topline model in certain aspects, indicating the effectiveness of our streamable architecture for zero-shot VC. Additionally, C-StreamVoice exhibits inferior performance compared to the discrete version, which can contribute to the over-smoothness in speech generation Ren et al. (2022) and the mismatch between the ground truth and predicted features.
As illustrated in Table. 2, the RTF of the entire pipeline is below 1, which meets the real-time requirement. Consisting of chunk-waiting latency (80ms) and model inference latency, the overall pipeline latency is 124.3 ms. If using a V100 GPU, StreamVoice can obtain an RTF of 0.56, and the overall latency reaches 137.2 ms. Importantly, unlike previous streaming VC, our VC model is entirely causal without any future look-ahead, highlighting its powerful modeling capability. These results show that StreamVoice can achieve high-quality zero-shot VC in streaming scenarios.
@@
Table 2: Speed tested on an A100 80G GPU. Latency is obtained by multiplying RTF by 80ms chunk size.
Method Quality Similarity
GT (origin) - 3.65 6.29 - 0.729
Non-streaming Topline
NS-VC 3.85$\pm$0.09 3.71 8.39 3.92$\pm$0.08 0.744
Streaming Model
IBF-VC 3.71$\pm$0.09 3.48 9.52 3.67$\pm$0.10 0.687
DualVC2 3.80$\pm$0.10 3.57 10.2 3.85$\pm$0.09 0.703
StreamVoice 3.82$\pm$0.09 3.50 10.0 3.82$\pm$0.10 0.694
$\quad$++ Tuning 3.78$\pm$0.08 3.52 10.4 3.87$\pm$0.10 0.714
Table 3: In-dataset performance (seen speakers)
@@ 4.2.2 In-dataset Evaluation
To get further insight into StreamVoice, we conducted an in-dataset evaluation on eight seen speakers, as shown in Table. 3. A non-streaming VC system Tian et al. (2020) achieving any2many VC, is selected, referred to as NS-VC. Also, IBF-VC Chen et al. (2022) and DualVC2 Ning et al. (2024) are recently proposed streaming models for any2many VC. As observed, a performance gap exists between the strong non-streaming topline and streaming models. Among the streaming models, StreamVoice, designed for the zero-shot scenario, delivers similar results to systems designed for in-dataset speakers, even though StreamVoice uses a smaller chunk size of 80ms in streaming ASR, achieving lower ASR performance. In contrast, IBF-VC and DualVC2 employ 160ms chunk size of ASR for streaming VC. It indicates StreamVoice’s good conversion ability. With available utterances of target speakers, fine-tuning yields superior performance. This indicates our system can be easily applied to various scenarios with or without the utterances of target speakers.
@@ 4.3 Ablation Study
As presented in Table 4, we conducted several ablations studies. In w/o teacher-guided context foresight, we discard the prediction of current and future semantic information, forming two ablations w/o $\mathcal{L}_{TF}(\overline{s}_{t})$ and w/o $\mathcal{L}_{TF}(\overline{s}_{t+1:t+k})$. As can be seen, a noticeable decrease occurs in all evaluation metrics when the $\mathcal{L}_{TF}(\overline{s}_{t+1:t+k})$ is discarded, especially in WVMOS and CER. This indicates that this foresight improves performance in capturing the linguistic content. But when only integrating context from future semantics, the model w/o $\mathcal{L}_{TF}(\overline{s}_{t})$ faces severe performance loss. It shows that only using future information interferes with delivering current linguistic content. In w/o semantic masking, we observe a performance decrease in all evaluation metrics when the semantic masking is discarded. This indicates that StreamVoice, trained with semantic masking, effectively enhances contextual learning from the preceding input while improving speaker timbre capturing. Furthermore, the results of dropping the bottleneck regulator show that its integration effectively prevents the source speaker information contained in the semantic feature from leaking into the converted speech, with little effect on speech quality.
@@ 4.4 Discussion: Dependency Analysis
In this section, we will explore the dependency relations between the selection of ASR and codec and the performance of StreamVoice.
ASR. To investigate the impact of ASR on StreamVoice, three representative ASR systems, including non-streaming ASR 4, widely used CTC-based streaming ASR Moritz et al. (2019), and the recently proposed streaming Fast-U2++ Liang et al. (2023), are selected to perform semantic extraction. As can be seen in Table 5, StreamVoice using semantic features of non-streaming ASR outperforms those using streaming ASR. This discrepancy may be attributed to the inherent performance gap between non-streaming and streaming ASR models, resulting in different semantic extraction abilities. Besides, without future look-ahead in StreamVoice, using semantic features from Moritz et al. (2019) cannot achieve reasonable conversion, while we introduce a 160ms future look-ahead in StreamVoice, i.e., modeling $p(a_{t}|\mathbf{a}_{1:t-1},\mathbf{s}_{1:t+m},t)$ with $m$ future look-ahead, yield good conversion results. This issue may arise from delayed CTC spike distributions and token emission latency existing in streaming ASR Liang et al. (2023), leading to semantic information shifting. Benefiting from the low emission latency of Fast-U2++, StreamVoice can perform conversion without future look-ahead. With a longer chunk size employed in Fast-U2++, StreamVoice can obtain better results while reaching a larger latency of 270ms. A trade-off still exists between performance and speed.
@@
Table 6: Analysis of dependency on Audiodec with various bitrate.
Codec. In StreamVoice, we employ a low-latency streaming codec Audiodec Wu et al. (2023). As presented in Table. 6, we validate the performance of StreamVoice using codecs with different bitrates, including 2kbps and 8kbps, where higher bitrate codecs achieve superior reconstruction quality to lower bitrate ones. The 2kbps Audiodec utilizes 4 layers of quantization and represents audio with a frame length of 20ms, while the 8kbps Audiodec employs 8 layers with a frame length of 10ms. Using the configuration of StreamVoice mentioned in Section 4.1, the results in different bitrates of codec models show no obvious differences. When increasing the number of transformer layers in the codec predictor, forming Large w/ 8kbps Audiodec, conversion performance using 8kbps codec improves noticeably, but resulting in slower inference. This result shows that the design of StreamVoice depends on the codec configuration, affecting both conversion quality and inference speed.
5 Conclusions
This paper introduces StreamVoice, a novel LM-based zero-shot VC system designed for streaming scenarios. Specifically, StreamVoice employs a single-stage framework encompassing a context-aware LM and an acoustic predictor. The casual design of the model’s input and structure ensures compliance with streaming behavior. To address performance degradation caused by missing complete contextual information in streaming scenarios, context-aware LM adopts teacher-guided context foresight to make the model able to forecast the current and future information given by a teacher. Besides, semantic masking is introduced in LM to enhance context learning from historical input and facilitate better disentanglement. Finally, an acoustic predictor collaborates with the LM to generate the target speech. Experiments demonstrate that StreamVoice achieves streaming zero-shot VC while maintaining performance comparable to non-streaming VC systems.
6 Limitations
We have to point out that StreamVoice still has limitations. In our configuration, StreamVoice needs a GPU, such as V100 and A100, to achieve real-time streaming inference. The design of streaming VC heavily relies on the ASR and the speech codec as mentioned in Section 4.4. Besides, StreamVoice also faces the out-of-domain problem, which causes performance degradation for utterances with accents, strong emotions, or unseen recording environments. Our future work will first use more training data with diversity coverage to explore StreamVoice’s modeling ability. Also, we will focus on optimizing our streaming pipeline, such as high-fidelity codec with low bitrate and unified streaming model.
7 Ethics Statement